OpenStudy (anonymous):

The length of a rectangle is 3 centimeters less than twice its width its area is 20 square centimeter. Find the dimensions of the rectangle. use the formula, area=length*width.

OpenStudy (zpupster):

write down what you know. area is 20 m2 20 = width we dont kno so let that be w our length 3 less than 2 times the width -3 + 2w =2w-3 our equation 20 = (w)(2w-3) 20 = 2w^2-3w 2w^2- 3w - 20 = 0 2w^2 - 8w - 5w - 20 = 0 2w(w-4)+5(w-4) = 0 (2w+5)(w-4)=0 w-4=0 w=4 l=2w-3 l=8-3 l=5 check 20 = 5(4)

OpenStudy (anonymous):

Thank you so much!!!! =]